21 kg mass is located on the ground the static friction coefficient is 0.75 and the kinetic friction coefficient is 0.65 the mass pushes with 215 N of force an angle of 20° below the horizontal


1
Expert's answer
2021-03-18T14:02:18-0400

GIVEN:

A block of 21 kg mass is placed onthe rough surface with

static friction coefficient "(\mu_s )=0.75"

kinetic friction coefficient "(\mu_k )=0.65"

applied force of 215 N at an angle "20\degree" below the horizontal


DIAGRAM :


(fig.1) shows how force is applied on the block.


Now, we can make component of fore in x and y directions

as F1 and F2 respectively.

"\implies" refer (fig. 2)



Values of component forces are as follows:

as, F=215 N and angle is "20\degree"

"\therefore F_1=215 cos(20\degree)
 =212.0339"

"F_2=215 Sin(20\degree)
 =73.5343"

"\implies refer (fig.3)"





now, normal force will increase and there is a +ve force along x direction.

Normal force

"N=mg +F_2"

"= (210+73.5343)"

"=283.5343"

"\implies refer(fig.4)"




When horizontal force F1 will work in x direction then box try to move and frictional force will come in action...

"\implies" so first we find maximum friction force "(f_m)" that can be applied by the box

"\boxed{f_m=f_s=\mu_sN}"

"=0.75*283.5343"

"\boxed{f_m=212.6507}" ....................(1)


"\implies" and force applied in x-deirection is

"\boxed{F_1=202.0339}" ................................(2)



"\bigstar" from (1) and (2) ,we get to know that maximum frictional force is more than horizontal force

"\therefore" frictional force "f" will be equal to "F_1"

i.e, 202.0399


"\bigstar" it implies that the box will not move and remain static on its position.